Evaluate the limit "\displaystyle{\lim\limits_{\theta \to 0} \dfrac{\sin{(\theta^2)}}{\theta}}" using the l'Hopital's Rule.


1
Expert's answer
2021-09-20T17:29:28-0400

"\lim\limits_{\theta\to0}{sin(\theta^2)\over \theta}"

"=\lim\limits_{\theta\to0}({f'(sin(\theta^2))\over f'( \theta)})"

"=\lim\limits_{\theta\to0}({2.\theta.cos(\theta^2)\over 1})"

"=({2.0.cos(0^2)\over 1})={0\over 1}=0"